OpenStudy (anonymous):

identify one of the two solutions: x^2+6x-40=0

OpenStudy (anonymous):

x^2 + 6x - 40 = 0 Lets find the roots of the equation first which can be 2,20 ad 4,10 and 5,8 But 4,10 will satisfy the equation since (10 -4 = 6) x^2 + 10x - 4x - 40 = 0 x (x + 10) -4 (x + 10) = 0 [Taking x common] (x-4) (x + 10) = 0 X - 4 = 0 => x = 4 x + 10 = 0 => x = - 10 So two solutions are x = 4 nd x = -10

OpenStudy (anonymous):

so if it says identify one do i just put one of those you gave me ?

OpenStudy (anonymous):

Since Root cannot be negative it should be the positive one so ur answer should be x = 4

OpenStudy (anonymous):

thanks a lot dude

OpenStudy (anonymous):

Anytime !

OpenStudy (anonymous):

i have one more lol if u dont mind i have an assingment due at night

OpenStudy (anonymous):

Sure np

OpenStudy (anonymous):

its the same thing but 4x^2+19x-5=0

OpenStudy (anonymous):

Roots would be 20 and 1 which satisfies 20 - 1 = 19 so it would be 4x^2 + 20x - x - 5 = 0 4x (x + 5) -1 (x + 5) = 0 (4x -1) (x + 5) = 0 4x - 1 = 0 => x = 1/4 x + 5 = 0 => x = -5 So if u want to pick one it would be 1/4

OpenStudy (anonymous):

so 1/4 is the answer.